What is the total electric flux through a closed surface containing a 2.0 μC charge?


1
Expert's answer
2021-04-19T04:34:22-0400

"\Phi=\frac{q}{\epsilon_0}=\frac{2\cdot 10^{-6}}{8.85\cdot 10^{-12}}=2.26\cdot 10^5~\frac{N\cdot m^2}{C}"